This is what i submitted and am being told that the composition of one of my answers is incorrect. I don't know which one.
(f ° g)(x) and (g ° f)(x)
f(x)=√(x+1),g(x)=x+2

f(x+2)=√(x+2+1)
=√(x+3)
=x+3
g(√(x+1))=√(x+1)+2
=x+1+2
=x+3

Answer by stanbon(75887)

(f ° g)(x) and (g ° f)(x)
f(x)=√(x+1),g(x)=x+2
-----------------------------
f[g(x)] = f[x+2) = sqrt(x+2+1) = sqrt(x+3)
---------------------
g[f(x)] = g[sqrt(x+1)] = sqrt(x+1)+2
=======================================
Cheers,
Stan H.
==========